# Automating Automotive Software Development: A Synergy of Generative AI and Formal Methods

### v1

@@ Abstract
As the automotive industry shifts its focus toward software-defined vehicles, the need for faster and reliable software development continues to grow. However, traditional methods show their limitations. The rise of Generative Artificial Intelligence (GenAI), particularly Large Language Models (LLMs), introduces new opportunities to automate automotive software development tasks such as requirement analysis and code generation. However, due to the complexity of automotive systems, where software components must interact with each other seamlessly, challenges remain in software integration and system-level validation. In this paper, we propose to combine GenAI with model-driven engineering to automate automotive software development. Our approach uses LLMs to convert free-text requirements into event chain descriptions and to generate platform-independent software components that realize the required functionality. At the same time, formal models are created based on event chain descriptions to support system validation and the generation of integration code for integrating generated software components in the whole vehicle system through middleware. This approach increases development automation while enabling formal analysis to improve system reliability. As a proof of concept, we used GPT-4o to implement our method and tested it in the CARLA simulation environment with ROS2 middleware. We evaluated the system in a simple Autonomous Emergency Braking scenario.
@@ I Introduction
Over the past decade, the automotive industry has been undergoing a transformation from the traditional vehicles focused on mechanical systems to software-defined vehicles (SDVs). As the amount of software in vehicles increases, the complexity of automotive software development is growing rapidly. Modern cars are now estimated to be about five times more complex than they were just a few years ago [1]. Meanwhile, productivity in software development is increasing only slowly, which creates a widening gap between rising demands and the available development capacity.
@@
Recently, the emergence of Generative Artificial Intelligence (GenAI), particularly Large Language Models (LLMs), has provided new possibilities for automating tasks across various domains, thanks to their strong capabilities in text understanding and generation [2]. Their potential is now being explored in the automotive sector as well [3, 4]. LLMs show promise for tasks such as requirement analysis, documentation, and even code generation. However, several challenges limit their practical application in the automotive domain.
@@
Automotive systems are highly complex, involving numerous software and hardware components interacting with each other. These systems typically follow a layered software architecture, which promotes modularity and scalability through the use of middleware. Although LLMs can assist in generating small software modules, creating complete and dependable system-level software remains a significant challenge. Furthermore, integrating these LLM-generated components into existing vehicle systems adds another layer of complexity. In addition, automotive platforms often require strict guarantees related to safety, real-time performance, and reliability. These constraints are difficult to analyze using purely statistical AI models. In such cases, system validation often depends on formal methods and expert review.
@@
In this work, we aim to accelerate automotive software development by combining the generative capabilities of LLMs with the rigor of formal methods. In the automotive domain, system behavior is often described using the event chain concept [5], which breaks down high-level software functionality into smaller, well-defined processing steps with clearly specified inputs and outputs. Event chains are frequently used not only to guide software implementation but also to support system analysis [6]. Building on this concept, we introduce an event chain-driven development approach based on an agentic workflow that coordinates multiple specialized LLMs. Our process begins by using an LLM to extract functional software behaviors from natural language requirements. These behaviors are structured into event chains, where each step is described as a subcomponent, representing a modular unit of functionality. For each subcomponent, middleware-independent software is generated using a code-specific LLM. In parallel, a formal representation of the event chain, e.g., in Eclipse Modeling Framework (EMF) format, is derived via a model-specific LLM. The formal event chain model allows for early-stage system validation and facilitates post-analysis. At the same time, the formal model enables model-based code generation to integrate subcomponents into the overall runtime environment. The completed software is then deployed to the vehicle platform or a simulation environment for further testing and analysis.
We demonstrate the proposed approach in a basic Autonomous Emergency Braking (AEB) scenario within a ROS2-based simulation environment using the CARLA simulator [7]. As the capabilities of cutting-edge LLMs are evolving on a daily basis, we do not intend to benchmark and identify the most suitable LLM for each individual step. Instead, we aim to provide a showcase of the proposed method. Thus, the well-known, general-purpose LLM, OpenAI’s GPT-4o, is employed for event chain construction, code generation, and formal modeling. This demonstration highlights that the proposed method can significantly automate the automotive software development process while retaining the ability to perform further system analysis and validation.
@@ II Related Work
Numerous studies have investigated the use of LLMs to generate automotive-related code from textual descriptions. Abdalla et al. [8] investigated fine-tuning small open-source LLMs to generate low-level vehicle control functions in graphical programming languages for model-based development environments, such as MATLAB Simulink. The resulting models were subsequently used to generate C code via traditional model-based techniques. Patil et al.  [9] focused on generating industrial-grade C code for automotive embedded systems using LLMs. Their work supports both high-level and low-level textual specifications, including the formal ANSI/ISO C Specification Language, and integrates formal code verification tools to analyze the generated output. Nouri et al. [10] proposed an iterative approach for generating safety-critical vehicle functions. They combined Python code generation with the esmini simulation environment and predefined test cases. If the generated code failed the test, the LLM was prompted again with the test results and previous code to trigger a correction cycle. These works demonstrate that modern LLMs are capable of producing syntactically and semantically valid code for automotive applications. However, they primarily focus on generating isolated functional scripts and do not address the challenges of integrating multiple software components or how these components interact within a complete vehicle system.
In parallel, advanced model-based approaches have been widely used to support system-level integration and analysis in automotive domain. Vinoth Kannan [11] provided an overview of model-based methods in automotive systems, highlighting how they can improve quality and reduce development time through verification and automated code generation. Holtmann et al. [12] presented a comprehensive model-based development pipeline, starting from system architecture in SysML to software design and integration. They proposed automating the transformation from SysML models to AUTOSAR-compliant software models, which enables seamless software integration into the vehicle runtime. Our previous work [13] explored the use of formal models to represent vehicle software and hardware for analyzing resource allocation issues. We also utilized model-based code generation to automate deployment-related software. These studies confirm the strengths of model-based techniques in automotive development. However, a key challenge remains for the usage of model-based method is significant domain expertise required to construct and use formal models. Recently, several works [14, 15, 16] have explored how advanced LLMs can be leveraged to generate model components automatically. This opens new opportunities to combine generative AI and formal methods to automate the development of automotive software systems.
@@ III Agentic Workflow
In parallel, a formal representation of the event chain is constructed using a modeling-specific LLM (Section IV-B). This produces a system model instance conforming to a predefined event chain meta model. The formal model can capture both the event chain’s architectural structure and any relevant system properties. In addition, formal constraints will be extracted and formalized from the design decisions, specifications and requirements (mainly non-functional related). These formal models and constraints support early-stage system validation (Section V-B) and facilitates post-simulation analysis by allowing the integration of runtime metrics back into the model. The formal instance model also enables model-based code generation, producing the integration code that connects individual subcomponents into the vehicle runtime system (Section V-C). The advantage of model-driven integration lies in its analytical approach, which significantly reduces the uncertainty compared to purely statistical approaches driven by LLMs. As a result, integration can be performed in a deterministic way, allowing testing efforts to concentrate on validating the correctness of the generated software components and their compliance with the functionality defined in requirements, rather than resolving integration errors or interface mismatches.
@@ IV LLM-based Generation
[…]
@@ IV-A Event Chain Generation
The prompt used guide the LLM in creating the event chain is provided in Listing 1. We utilize a standardized JSON format to structure the relevant software and signal information. This includes component names, textual descriptions, implementation logic, and detailed input/output signal specifications. The JSON template for software components is presented in Listing 2, while the structure for signal descriptions is shown in Listing 3. As our prototype is implemented using ROS, the signal templates in our examples include ROS-specific fields such as topic names and message types. However, this approach is not limited to ROS. A more generalized signal definition, such as Vehicle Signal Specification (VSS), can also be adopted. This would require adapters during implementation to convert standardized signal formats into platform-specific implementations. However, the discussion of standardized signal representation and its integration with runtime middleware is beyond the scope of this paper.
@@
The resulting event chain is represented as a list of software components in JSON format. Each component defines its required input/output signal and its implementation logics based on the requirements. This structured event chain description is then passed on to the next stage for software code generation (Section IV-C) and instance model generation (Section IV-B).
Listing 1: Prompt for event chain generation

### v2

# Automating Automotive Software Development: A Synergy of Generative AI and Model-Based Methods

@@ Abstract
As the automotive industry shifts its focus toward software-defined vehicles, the need for faster and reliable software development continues to grow. However, traditional methods show their limitations. The rise of Generative Artificial Intelligence (GenAI), particularly Large Language Models (LLMs), introduces new opportunities to automate automotive software development tasks such as requirement analysis and code generation. However, due to the complexity of automotive systems, where software components must interact with each other seamlessly, challenges remain in software integration and system-level validation. In this paper, we propose to combine GenAI with model-driven engineering to automate automotive software development. Our approach uses LLMs to convert free-text requirements into event chain descriptions and to generate platform-independent software components that realize the required functionality. At the same time, formal models are created based on event chain descriptions to support system validation and the generation of integration code for integrating generated software components in the whole vehicle system through middleware. This approach increases development automation via GenAI while enabling model-based analysis to improve system reliability. We demonstrated our method with different LLMs and tested it in the CARLA simulation environment with ROS2 middleware. We evaluated the system in a simple Autonomous Emergency Braking scenario.
@@ I Introduction
Over the past decade, the automotive industry has been undergoing a transformation from the traditional vehicles focused on mechanical systems to software-defined vehicles (SDVs). As the amount of software in vehicles increases, the complexity of automotive software development is growing rapidly. Modern cars are now estimated to be about five times more complex than they were just a few years ago [mckinsey]. Meanwhile, productivity in software development is increasing only slowly, which creates a widening gap between rising demands and the available development capacity.
@@
Recently, the emergence of Generative Artificial Intelligence (GenAI), particularly Large Language Models (LLMs), has provided new possibilities for automating tasks across various domains, thanks to their strong capabilities in text understanding and generation [matarazzo2025survey]. Their potential is now being explored in the automotive sector as well [phatale2024, petrovic2024-3]. LLMs show promise for tasks such as requirement analysis, documentation, and even code generation. However, several challenges limit their practical application in the automotive domain.
@@
Automotive systems are highly complex, involving numerous software and hardware components interacting with each other. These systems typically follow a layered software architecture, which promotes modularity and scalability through the use of middleware. Although LLMs can assist in generating small software modules, creating complete and dependable system-level software remains a significant challenge. Furthermore, integrating these LLM-generated components into existing vehicle systems adds another layer of complexity. In addition, automotive platforms often require strict guarantees related to safety, real-time performance, and reliability. These constraints are difficult to analyze using purely statistical AI models. In such cases, system validation often depends on model-based formal methods and expert review.
@@
In this work, we aim to accelerate automotive software development by combining the generative capabilities of LLMs with the rigor of model-based methods. In the automotive domain, system behavior is often described using the event chain concept [muenzenberger2021event], which breaks down high-level software functionality into smaller, well-defined processing steps with clearly specified inputs and outputs. Event chains are frequently used not only to guide software implementation but also to support system analysis [padma2020]. Building on this concept, we introduce an event chain-driven development approach based in an agentic workflow that coordinates multiple specialized LLMs. Our process begins by using an LLM to extract functional software behaviors from natural language requirements. These behaviors are structured into event chains, where each step is described as a subcomponent, representing a modular unit of functionality. For each subcomponent, middleware-independent software is generated using a code-specific LLM. In parallel, a formal representation of the event chain, e.g., in Eclipse Modeling Framework (EMF) format, is derived via a model-specific LLM. The formal event chain model allows for early-stage system validation and facilitates post-analysis. At the same time, the formal model enables model-based code generation to integrate subcomponents into the overall runtime environment. The completed software is then deployed to the vehicle platform or a simulation environment for further testing and analysis.
We demonstrate the feasibility of the proposed approach through a basic Autonomous Emergency Braking (AEB) scenario within a ROS2-based simulation environment using the CARLA simulator [dosovitskiy2017carla]. Several LLMs, including GPT-4o, Gemini 2.5 Pro, and Llama 3.3 70B, are evaluated for their effectiveness in the proposed workflow.
@@ II Related Work
Numerous studies have investigated the use of LLMs to generate automotive-related code from textual descriptions. Abdalla et al. [abdalla2024] investigated fine-tuning small open-source LLMs to generate low-level vehicle control functions in graphical programming languages for model-based development environments, such as MATLAB Simulink. The resulting models were subsequently used to generate C code via traditional model-based techniques. Patil et al.  [patil2024] focused on generating industrial-grade C code for automotive embedded systems using LLMs. Their work supports both high-level and low-level textual specifications, including the formal ANSI/ISO C Specification Language, and integrates formal code verification tools to analyze the generated output. Nouri et al. [nouri2025] proposed an iterative approach for generating safety-critical vehicle functions. They combined Python code generation with the esmini simulation environment and predefined test cases. If the generated code failed the test, the LLM was prompted again with the test results and previous code to trigger a correction cycle. These works demonstrate that modern LLMs are capable of producing syntactically and semantically valid code for automotive applications. However, they primarily focus on generating isolated functional scripts and do not address the challenges of integrating multiple software components or how these components interact within a complete vehicle system.
In parallel, advanced model-based approaches have been widely used to support system-level integration and analysis in automotive domain. Vinoth Kannan [VinothKannan2021] provided an overview of model-based methods in automotive systems, highlighting how they can improve quality and reduce development time through verification and automated code generation. Holtmann et al. [holtmann2021] presented a comprehensive model-based development pipeline, starting from system architecture in SysML to software design and integration. They proposed automating the transformation from SysML models to AUTOSAR-compliant software models, which enables seamless software integration into the vehicle runtime. Our previous work [pan2024sdv] explored the use of formal models to represent vehicle software and hardware for analyzing resource allocation issues. We also utilized model-based code generation to automate deployment-related software. These studies confirm the strengths of model-based techniques in automotive development. However, a key challenge remains for the usage of model-based method is significant domain expertise required to construct and use formal models. Recently, several works [petrovic2025meta, pan2025model, pan2024ocl] have explored how advanced LLMs can be leveraged to generate model components automatically. This opens new opportunities to combine generative AI and formal methods to automate the development of automotive software systems.
@@ III Agentic Workflow
In parallel, a formal representation of the event chain is constructed using a modeling-specific LLM (Section IV-B). This produces a system model instance conforming to a predefined event chain meta model. The formal model can capture both the event chain’s architectural structure and any relevant system properties. In addition, formal constraints will be extracted and formalized from the design decisions, specifications, and requirements (mainly non-functional) manually or through LLMs. These formal models and constraints support early-stage system validation (Section V-B) and facilitates post-simulation analysis by allowing the integration of runtime metrics back into the model. The formal instance model also enables model-based code generation, producing the integration code that connects individual subcomponents into the vehicle runtime system (Section V-C). The advantage of model-driven integration lies in its analytical approach, which significantly reduces the uncertainty compared to purely statistical approaches driven by LLMs. As a result, integration can be performed in a deterministic way, allowing testing efforts to concentrate on validating the correctness of the generated software components and their compliance with the functionality defined in requirements, rather than resolving integration errors or interface mismatches.
@@ IV LLM-based Generation
The proposed workflow leverages Large Language Models at multiple stages for both requirement analysis and artifact generation. In the first step, an LLM is used to derive a structured event chain description from the provided requirements. Subsequently, the event chain description is modeled into an instance model by an LLM, which then serves as a crucial input for model-based analysis and further development. In addition, LLMs are also employed to generate software components that are not available in the existing software base.
@@ IV-A Event Chain Generation
The prompt used to guide the LLM in creating the event chain is provided in Appendix A. We utilize a standardized JSON format to structure the relevant software and signal information. This includes component names, textual descriptions, implementation logic, and detailed input/output signal specifications. As our prototype is implemented using ROS, the signal templates in our examples include ROS-specific fields such as topic names and message types. However, this approach is not limited to ROS. A more generalized signal definition, such as Vehicle Signal Specification (VSS), can also be adopted. This would require adapters during implementation to convert standardized signal formats into platform-specific implementations. However, the discussion of standardized signal representation and its integration with runtime middleware is beyond the scope of this paper.
@@
The resulting event chain is represented as a list of software components in JSON format. Each component defines its required input/output signal and its implementation logic based on the requirements. This structured event chain description is then passed on to the next stage for software code generation (Section IV-C) and instance model generation (Section IV-B).
IV-B Model and Constraints Creation
We follow the model generation approach proposed in [pan2025model] to generate a formal event chain model. This method leverages the capabilities of LLMs to generate a conceptual instance model in JSON format, which is subsequently parsed into a valid instance model file.
This step takes as input a self-defined event chain meta model (Section V-A) along with the event chain description produced from the previous step (Section IV-A). In our implementation, the event chain description is provided in a customized JSON format. Following the strategy outlined in [pan2025model], we manually construct a one-shot example comprising the event chain meta model, a similar event chain description, and the expected instance model. This example is used to guide the LLM in generating the conceptual instance model for the target system. The output from the LLM is then parsed by the instance modeler introduced in [pan2025model], resulting in a structured instance model that conforms to the predefined meta model. This model serves as a foundation for subsequent tasks, including model-based system analysis and model-driven code generation (Section V-C).
The constraint generation is a relatively straightforward process, as mentioned in [pan2024ocl]. In our case study, we employ the Object Constraint Language (OCL) [OMG2014] as the constraint language for model-based formal constraint description. It has logic similar to first-order logic and is designed for MOF-based model validation. We simply pass the meta model information and natural language text to the LLM, and the corresponding OCL constraints are generated. Examples of generated OCL constraints are presented in Section V-B
IV-C Code generation
We consider each subcomponent described in the event chain as a standalone piece of software that can be developed and tested independently before integration into the complete system. The event chain description of a subcomponent includes its behavior and implementation logic, both derived from the system requirements. This information is provided to the LLM through a structured prompt to guide the generation of code for the subcomponents. An example of such a prompt is shown in Appendix B.
V Model-based Formal Approach
V-A Meta model and instance model
Models are fundamental elements in any model-based methodology. According to the modeling standard, the Object Management Group’s Meta-Object Facility (OMG MOF) [OMG2016], models are categorized to meta models and instance models based on their level of abstraction. The meta model defines the abstract information of a group of systems. It serves as the modeling syntax or schema for constructing corresponding instance models. In contrast, an instance model captures detailed information about a specific system and adheres to the structure defined by its corresponding meta model. Widely adopted modeling languages and tools such as Unified Modeling Language [pilone2005uml], Systems Modeling Language [friedenthal2014practical] and Eclipse Modeling Framework [Steinberg2008] are compatible to this standard.
To formally model event chains, we developed an example meta-model using EMF, as illustrated in Figure 2. In our meta model, an event chain may consist of multiple software components and associated data elements. These data elements represent input and output signals of the software components. Attributes such as execution frequency, data topics, and message types are specified. These details are essential for enabling model-based code generation. Additional attributes can also be defined in the meta model for enabling further system analysis. However, comprehensive modeling strategies for enabling model-based system analysis will not be further discussed in this paper.
In this work, we employ LLMs to automatically generate instance models based on our defined meta-model (Section IV-B). An example of such a generated instance model, depicting a basic AEB system , is shown in Figure 3. This AEB system serves as case study scenario in Section VI.
Figure 2: An example meta model for event chain
V-B Constraints and Model Validation
context SoftwareNode
inv HasInputAndOutputData:
self.input->notEmpty() and self.output->notEmpty()
inv NextstepFrequencyEqualOrHigher:
self.nextstep->notEmpty() implies (self.nextstep.
frequency >= self.frequency)
Using the Eclipse OCL Plugin or other solver-based formal methods mentioned in [pan2024sdv], model validation can be performed and unsatisfied constraints with respect to a specific instance model can be detected.
V-C Model-based Code Generation
We employ model-to-text generation techniques to produce middleware-specific code that integrates individual subcomponents into the vehicle runtime. For models adhering to the OMG MOF standard, the OMG MOF Model to Text Transformation Language (MOFM2T) specification defines a standard method for generating textual artifacts from MOF-based models. We use Acceleo [acceleo], a model-to-text generation tool, to generate code from the information contained within instance models.
An example code template for generating ROS-specific code to integrate LLM-generated software components in our demonstration scenario is presented in Appendix C. The red text indicates AQL operations, including the entry point of generation. The green text represents comments, which also provide clarification and refer to the main Acceleo class implementation. The blue text denotes AQL queries. These queries are written using meta-information (e.g., class and property definitions) from the meta model and are used to retrieve specific objects defined in the instance model. The rest of texts are predefined static codes constructing the class of a ROS node. The initial lines of the template specify generation configuration, including the entry point, which in our case is the EventChain class. We then iterate over each SoftwareNode within the event chain and retrieve required attributes through AQL queries, generating a corresponding Python script for each as an individual ROS node that can be directly deployed into ROS runtime. The node structure is predefined within the template. Signal publishing, subscription mechanisms, and submodule interactions are generated based on the data preserved in the instance model. In general, each generated ROS node receives signals, passes them to the corresponding LLM-generated core module, obtains the result, and then publishes it for subsequent processing.
Figure 3: Generated instance model for the AEB case study
VI Experiment
In this section, we demonstrate the feasibility of the proposed automotive software development approach using a basic AEB scenario.
In the implementation, we utilized Eclipse IDE 2022-06 to create the event chain meta-model (Fig. 2) and to present the generated instance model (Fig. 3). The OCL All-In-One SDK (version 6.17.1.v20220309-0840) was used for model validation. The Acceleo tool (version 3.7.11.202102190929) served as the engine for model-to-text code generation.
As LLMs, we used and compared GPT-4o, Gemini 2.5 Pro and Llama 3.3 70B for their performance in generating models and codes. The commercial LLMs, GPT-4o and Gemini 2.5 Pro, were accessed via their official APIs. The open source Llama 3.3 70B was executed through the Groq platform. The experiment of each LLM was conducted 5 times. Each LLM was evaluated across five independent runs. The comparison results are presented in Section VI-B.
We further set up a CARLA simulation to verify the generated code. We utilize CARLA version 0.9.15, and choose Town 01 as the map. A straight lane before the traffic light replicates a typical scenario that could trigger the brake system. A stopped vehicle is spawned before the red light, and our ego vehicle should stop before crashing. A Carla-ros-bridge has been built to exchange sensor and actuator information between ROS and the CARLA environment. Specifically, the ROS2 Foxy version of the bridge is used for improved real-time performance. The simulation runs on a PC equipped with an Intel Core i7-6850K CPU, 32 GB RAM, and two Nvidia GTX 1080 GPUs.
VI-A AEB Case Study
This subsection presents one successful end-to-end execution using GPT-4o. The following requirements were used for generating the AEB software:
  • The system shall determine brake force based on TTC thresholds: Full brake if TTC << 1.0s, Partial brake if 1.0s $\leq$ TTC << 2.0s, No brake if TTC $\geq$ 2.0s
  • The system shall output a normalized brake force (0.0 to 1.0).
As a preliminary step, we constructed two ROS nodes to serve as existing software components. The ObjectDetection node integrates the open-source pointcloud_to_laserscan package [pointcloudtolaser2025], which converts LiDAR point cloud data from CARLA into LaserScan messages containing the shortest distances to nearby obstacles. The VehicleControl node receives control commands and translates them into vehicle control messages for braking and other maneuvers. Additionally, two dummy software components were included for simulation purposes. Input information about the existing ROS topics and their message definitions was also collected.
Figure 4: Illustration of generated AEB event chain
By comparing the event chain description with the list of existing components, we identified that TTC_Calculation and BrakingDecision were not yet implemented and needed to be generated. We passed their descriptions to the prompt in Appendix B and generated the function code as individual Python classes for each component.
Figure 5: Functioning in CARLA simulator
Simultaneously, the entire event chain was modeled as a formal EMF instance (Fig. 3) based on the event chain meta-model (Fig. 2) for further inspection and validation. The instance model successfully passed the early-phase validation of the OCL constraints presented in Section V-B using Eclipse OCL. Afterwards, the instance model was utilized via model-based code generation (Section V-C) to generate integration-related code that integrates the function code of each component as ROS nodes.
The AEB software was executed in the predefined simulation scenario. The generated AEB module was successfully triggered, and the ego vehicle responded by braking in time before reaching the stationary vehicle (Fig. 5). This basic AEB system demonstrates the feasibility of our development approach.
VI-B Comparison of LLMs
To evaluate the performance of different LLMs in our agentic workflow, we measured their ability to generate both event chain models and functional code components. Each LLM was tested in five complete development runs.
We expect the generated artifacts to be accurate both semantically (with respect to the target modeling and programming languages) and syntactically (by conforming to the given system requirements). The evaluation results are summarized in Table I. To be considered successful, a single execution must result in both a valid event chain model and functional code, and the resulting AEB system must perform correctly within the simulation environment.
Among the evaluated models, GPT-4o achieved the highest overall success rate (60%), showing strongest performance in event chain modeling (80%). Gemini 2.5 Pro produced the most reliable function code (100%) but exhibited lower overall performance due to difficulties in generating valid event chain models. In contrast, Llama 3.3 70B showed limited effectiveness in both tasks, with no successful end-to-end executions.
These results highlight the complementary strengths of different LLMs, suggesting that a multi-agent workflow, where distinct LLMs are assigned to specific sub-tasks, could enhance the robustness and reliability of the overall development pipeline.
TABLE I: Accuracy of generated artifacts
LLM eventchain model function code overall success
GPT-4o 80% 60% 60%
Gemini 2.5 Pro 40% 100% 40%
Llama 3.3 70B 20% 20% 0%
VI-C Further discussion
The proposed approach combines model-based engineering with generative AI to support the development of automotive software systems. The model-based component ensures reliable system integration by addressing non-functional aspects such as safety requirements and design constraints. Meanwhile, LLMs are employed to enhance automation in development activities, including functional requirement analysis, code generation, etc. Due to the statistical nature of LLMs, the quality correctness of generated functional behavior cannot be guaranteed. Rigorous testing remains essential to validate the generated artifacts.
In conventional software development, testing, including unit testing, integration testing, and system testing, is critical for ensuring software quality. Within our workflow, integration code is generated using a correctness-by-construction approach, leveraging formal model definitions and predefined code templates. This eliminates the need for manual integration testing, as structural correctness is inherently assured by the model-based code generation process. However, because the functional code is generated by LLMs, its behavior must still be validated through dedicated testing procedures. Consequently, in the proposed workflow, testing efforts can be concentrated on unit testing of LLM-generated components and system-level testing of the overall vehicle behavior.
VII Conclusion
In this work, we introduced an agentic approach that combines the emerging capabilities of generative AI with model-driven formal methods to automate the development of automotive software. The approach leverages Large Language Models to analyze requirements, construct the overall software design as an event chain model, and generate platform-independent software function code as standalone classes. The formal event chain model serves as a basis for further (automated or human-centered) model-based system analysis and is also used to generate integration code, enabling the seamless connection of independent software components to the broader vehicle system via middleware. The resulting system can be deployed and evaluated in a simulation environment, such as CARLA. Simulation data can then be fed back into the event chain model to support further analysis, particularly in relation to non-functional requirements. We demonstrated the feasibility of this approach through a basic AEB scenario.
Acknowledgment
This research was funded by the Federal Ministry of Research, Technology and Space of Germany as part of the CeCaS project, FKZ: 16ME0800K.
Appendix A Prompt for event chain generation